### v1

# Consistency-Aware Parameter-Preserving Knowledge Editing Framework for Multi-Hop Question Answering

Abstract
Parameter-Preserving Knowledge Editing (PPKE) enables updating models with new or corrected information without retraining or parameter adjustment. Recent PPKE approaches based on knowledge graphs (KG) to extend knowledge editing (KE) capabilities to multi-hop question answering (MHQA). However, these methods often lack consistency, leading to knowledge contamination, unstable updates, and retrieval behaviors that fail to reflect the intended edits. Such inconsistencies undermine the reliability of PPKE in multi-hop reasoning. We present CAPE-KG, Consistency-Aware Parameter-Preserving Editing with Knowledge Graphs, a novel consistency-aware framework for PPKE on MHQA. CAPE-KG ensures KG construction, update, and retrieval are always aligned with the requirements of the MHQA task, maintaining coherent reasoning over both unedited and edited knowledge. Extensive experiments on the MQuAKE benchmark show accuracy improvements in PPKE performance for MHQA, demonstrating the effectiveness of addressing consistency in PPKE. All the codes are available at github.
Index Terms—  Knowledge Editing, Multi-Hop Question Answering, Knowledge Graph, Consistency, Parameter-Preserving Knowledge Editing
1 Introduction
00footnotetext: Corresponding author.
Knowledge editing [1] enables large language models (LLMs) to incorporate new or corrected information without the cost of full retraining. KE approaches can be divided into parameter-updating methods and parameter-preserving methods [2]. Parameter-updating methods, such as ROME [3], MEMIT [4], and T-Patcher [5], directly modify model weights to encode edits, which are computationally expensive. In contrast, PPKE avoids changing parameters and instead relies on external structures or contexts, offering greater transparency and stability. IKE [6] edits knowledge through in-context demonstrations, MeLLo [4] supports multi-hop editing through query decomposition and fact retrieval, and PokeMQA [7] performs decomposition and conflict checking. PPKE offers a cheaper and more efficient paradigm by avoiding both retraining and parameter modification. While early efforts in PPKE primarily addressed simple single-hop factual updates [6] [8], research has extended its scope to multi-hop question answering (MHQA) [9] [10], a more challenging setting that requires models to integrate edited knowledge with existing facts through multi-step reasoning. Recent work [11] adopts KG-based PPKE, utilizing KGs’ structured nature to enhance PPKE’s effectiveness in MHQA. By injecting edits into a graph and retrieving paths over entities and relations, Kedkg [11] allows models to perform explicit reasoning under edits.
However, current KG-based PPKE [11] is facing a critical weakness: lacking consistency [10]. Specifically, it struggles to maintain a knowledge boundary between factual and edited knowledge, to guarantee stable and conflict-free updates, and to ensure retrieval remains aligned with the intent of edit [12]. Inconsistency across these dimensions leads to less effective and reliable KE [13]. In this paper, we propose CAPE-KG, Consistency-Aware Parameter-Preserving Editing with Knowledge Graphs, the first consistency-aware framework for PPKE. It improves consistency across the entire pipeline, ensuring that KG construction, knowledge update, and knowledge retrieval are all aligned with the objective of the PPKE task. Our main contributions are:
We introduce a multi-layer KG architecture ensuring the consistent knowledge boundary between factual and edited knowledge during PPKE.
We design case-isolated updating mechanisms with conflict arbitration that guarantee edits are applied atomically and align with the case-isolated assumption.
We present an edit-aware retrieval module with progressive strategies and edit-related retrieval routing, ensuring the retrieval process is always consistent with the intended KE.
Extensive experiments on the MQuAKE benchmark [14] demonstrate that CAPE-KG consistently outperforms existing baselines, achieving substantial gains in consistency and accuracy across backbones.
2 Problem Formulation
Let $\mathcal{F}=\{(s,r,o)\}$ denote the set of factual triples, where $s$ is a subject, $r$ a relation, and $o$ an object; a knowledge edit is defined as $e=(s,r,o\rightarrow o^{*})$, specifying that the mapping $(s,r)\mapsto o$ should be updated to a new object $o^{*}$, and $\mathcal{E}=\{e_{1},e_{2},\ldots,e_{m}\}$ denotes the set of such edits. PPKE adds the constraint of avoiding adjustments to the model’s original parameters. For MHQA under PPKE, a multi-hop question $Q$ requires reasoning over a chain of facts $C=\big[(s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}),\ldots,(s_{n},r_{n},o_{n})\big]$, where $s_{i+1}=o_{i}$ and the final answer is $o_{n}$; edits in $\mathcal{E}$ revise relevant triples in $C$ to form $C^{*}=\big[(s_{1},r_{1},o_{1}^{\prime}),(s_{2},r_{2},o_{2}^{\prime}),\ldots,(s_{n},r_{n},o_{n}^{\prime})\big]$, with $o_{i}^{\prime}=o_{i}^{*}$. The objective is given $Q$, $\mathcal{F}$, and $\mathcal{E}$, the PPKE method would output an answer consistent with $C^{*}$.
Assumptions We adopt the case-isolated edit assumption, under which each edit is restricted to the current instance and does not influence other cases. This assumption aligns with the evaluation setting and reflects practical scenarios where distinct users or domains require isolated edits that must not interfere with unrelated reasoning chains.
3 Method
Fig. 1: An overview of CAPE-KG, which consists of three major components: the KG Construction module to build a pre-edit KG, the Knowledge Update module to update the KG according to edited knowledge, and the Knowledge Retrieval module to generate an answer through retrieval.
We propose the first consistency-aware PPKE framework, CAPE-KG, which addresses the fundamental challenges of three key dimensions of consistency. Fig.1 illustrates the overall pipeline, which consists of three major components: Knowledge Update, Knowledge Graph Construction, and Knowledge Retrieval. Each primarily ensures one form of consistency while reinforcing the others, resulting in a coherent editing framework. CAPE-KG maintains knowledge boundary consistency, where factual and edited knowledge is uncontaminated for reasoning; update consistency, where edits are incorporated in a nonconflicting and case-isolated manner; and intent consistency, where the whole retrieving process is always aligned with the purpose of the edit.
3.1 Knowledge Graph Construction
The construction of KG establishes the knowledge boundary consistency foundation by explicitly separating factual knowledge from edited knowledge. Prior Approach [11] constructs a single KG of edited knowledge, which creates knowledge boundary inconsistency: factual queries may fall back to LLM responses prone to hallucination or mistakenly retrieve edited knowledge from unrelated cases as facts. CAPE-KG adopts a multi-layer architecture, consisting of a base layer and corresponding overlay layers.
Base Layer. The base KG $B$ is constructed solely from factual triples $\mathcal{F}$, derived from the knowledge base. By excluding all edited content, this layer establishes a consistent foundation for factual reasoning.
Overlay Layers. For each editing case $c$, we construct an overlay $O_{c}=\{(s,r,o\to o^{*})\}$ containing only the edits associated with that case. Overlays are case-specific edit containers built on demand using a copy-on-write mechanism.
Knowledge Boundary Isolation. The separation between $B$ and $\{O_{c}\}$ guarantees that factual and edited knowledge remain distinct with respect to the knowledge boundary. This KG construction prevents edited knowledge from being misinterpreted as factual information, thereby maintaining knowledge boundary consistency across all reasoning processes. It provides the foundation for knowledge boundary consistency, which is further reinforced during each update and retrieval.
Method MQUAKE-CF-3K [14] MQUAKE-T [14]
I edited 100 edited All edited I edited All edited
M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc
LLaMa 2-7B [15]
FTCOT[16] 22.30 - 2.13 - OOM - 47.32 - 3.75 -
FT* [16] 28.20 7.30 2.37 0.03 OOM OOM 56.48 33.89 1.02 0.37
ROMECOT* 11.17 - 2.87 - 2.77 - 28.96 - 14.40 -
ROME* [3] 13.13 5.37 3.50 0.03 3.63 0.10 24.89 17.99 1.71 0.32
MEMITCOT[4] 11.83 - 9.23 - 5.57 - 36.88 - 31.58 -
MEMIT* [4] 14.97 6.43 9.40 2.47 2.30 0.37 30.89 23.98 25.21 20.13
MeLLo* [14] 33.57 9.90 20.00 10.07 17.33 9.90 65.78 55.27 57.69 44.55
PokeMQA* [7] 44.13 30.90 37.33 27.83 32.83 23.87 75.43 60.44 74.36 60.22
KeDKG* [11] 66.80 63.67 58.50 55.37 48.30 43.90 73.13 69.06 71.15 66.76
CAPE-KG 78.03 70.35 78.93 71.02 78.38 70.84 93.39 84.52 92.57 83.82
Vicuna-7B [17]
MeLLo* [14] 30.70 20.84 24.75 12.25 22.35 10.85 60.72 48.55 51.55 42.97
PokeMQA* [7] 45.83 34.80 38.77 31.23 31.63 25.30 74.57 55.19 73.07 55.09
KeDKG* [11] 68.60 65.13 62.43 58.20 51.10 44.67 71.90 67.23 74.68 66.64
CAPE-KG 81.89 78.74 81.59 78.85 80.94 78.49 98.08 89.14 97.45 88.20
GPT-3.5-turbo-instruct
MeLLo* [14] 57.43 28.80 40.87 28.13 35.27 25.30 88.12 52.84 74.57 53.53
PokeMQA* [7] 67.27 56.37 56.00 49.63 45.87 39.77 76.98 68.09 78.16 67.88
KeDKG* [11] 68.00 65.33 59.50 56.80 49.10 43.17 78.75 76.18 77.19 73.77
CAPE-KG 79.78 73.65 79.81 73.75 79.03 73.00 97.11 88.00 96.72 87.40
Table 1: The main experimental results across three backbone models: LLaMa2-7B [15], Vicuna-7B [17], and GPT-3.5-turbo-instruct. Metrics are M-Acc and H-Acc. Bold numbers indicate the best performance. Results marked with * are reported from KeDKG [11]. ”OOM” indicates out-of-memory failure due to GPU limits. ”k edited” is the number of edits within one batch.
3.2 Knowledge Update
Knowledge update ensures knowledge edits are incorporated in a stable and case-isolated manner. Given an input edit, we adopt the entity and relation detectors introduced in KEDKG [11] to extract a structured triple $\{(s,r,o^{*})\}$, which is then used to update the overlay of the current case. Unlike prior methods, where edits from different cases are repeatedly applied to a global graph [11], which leads to overwrites and unstable states, CAPE-KG enforces update consistency through a case isolation mechanism with conflict arbitration.
Updates are strictly limited to the overlay of the current case. By never propagating to the base layer or to overlays of other cases, edits remain aligned with the assumption that cases are independent. To enforce update consistency, the conflict arbitration module applies a deterministic rule: within each case, the current edit takes precedence for a given subject–relation pair; if no edit is specified, the original factual triple from the base layer is retained. This mechanism not only maintains update consistency but also reinforces the knowledge boundary separation established in construction, ensuring the boundary consistency throughout the system.
Fig. 2: Case Comparison of Update and KG Construction. Unedited knowledge is labeled in purple. Edited entities are in yellow, with dotted lines indicating the edited relation.
Fig. 3: Case Comparison of Intent Consistency in Retrieval.
Method MQUAKE-CF-3K [14] MQUAKE-CF-T [14]
KG Construction Retrieval Update I edited 100 edited All edited I edited All edited
M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc M-Acc H-Acc
75.00 69.00 74.67 69.04 75.67 69.00 96.00 86.60 95.80 86.60
53.33 43.00 51.33 46.67 51.93 46.00 87.33 81.67 87.67 82.00
77.40 71.00 74.58 68.14 74.29 68.10 96.60 87.10 94.35 85.60
79.78 73.65 79.81 73.75 79.03 73.00 97.11 88.00 96.72 87.40
Table 2: Ablation study of CAPE-KG on MQuAKE-CF-3K [14] and MQuAKE-T [14]. Each component incrementally contributes to improved consistency and accuracy, with the full framework achieving the best results.
3.3 Knowledge Retrieval
In the retrieval stage, an input query $Q$ is decomposed into a sequence of sub-questions $\{q_{1},q_{2},\ldots,q_{n}\}$ using a similarity-based dynamic few-shot strategy [18]. In Kedkg [11], each sub-question is answered by retrieving entities and relations from the knowledge graph. Existing methods, however, often fall back to unedited factual knowledge through LLM responses when retrieval fails, which can produce answers that contradict the intended edit. To address this, our framework designs a retrieval process that explicitly routes queries between the base and overlay layers, applies progressive strategies when uncertainty arises, and suppresses irrelevant edits.
Edit-Aware Retrieval Routing. We define the edit impact surface for case $c$ as $S_{\text{edit}}^{c}=\{s\mid(s,r,o^{*})\in E^{c}\}$ and $P_{\text{edit}}^{c}=\{r\mid(s,r,o^{*})\in E^{c}\}$, where $E^{c}$ represents the edit set for case $c$. For each sub-question $q_{i}$, we determine the target knowledge layer:
$\small L(q_{i})=\begin{cases}\text{Overlay}^{c},&\text{if subject}(q_{i})\in S_{\text{edit}}^{c}\lor\text{relation}(q_{i})\in P_{\text{edit}}^{c}\\ \text{Base},&\text{otherwise}\end{cases}$ (1)
This routing ensures that factual queries access the Base layer while edit-sensitive queries access the current case’s Overlay layer.
Progressive Retrieval. When answering a sub-question, the system may encounter uncertainty regarding entity or relation matching. To address this, we introduce a progressive retrieval strategy with three stages.
High Confidence Stage: For entity candidates $K=\{k_{1},\ldots,k_{n}\}$ with scores $g_{\phi}(q_{i},k_{j})\in[0,1]$ generated by detector from Kedkg [11], we first select those above a threshold $\tau$, $K^{\prime}={k_{j}\in K\mid g_{\phi}(q_{i},k_{j})\geq\tau}$. We then filter out low outliers using mean $\mu$ and deviation $\sigma$, keeping $K^{\prime\prime}={k_{j}\in K^{\prime}\mid g_{\phi}(q_{i},k_{j})\geq\mu-\lambda\sigma}$ where $\lambda>0$ is adaptive. Edit irrelevance suppression is performed during entity scoring, where confidence scores of edited entities not explicitly referenced in the sub-question are down-weighted to prevent unintended influence. The remaining candidates are sorted and sequentially used for retrieval, ensuring stable results when multiple entities have similar scores.
Low Confidence Stage: When no candidates exceed the threshold, i.e., $K^{\prime}=\emptyset$, we invoke a language model to select the most plausible entity from the original candidate pool $K$.
Failure Stage: If no valid entity is retrieved, a relation detector checks whether the subquestion $q_{i}$ falls within the scope of an edit. When relevant, the LLM is prompted with the full edited triple $(s,r,o^{*})$ injected into the context; otherwise, it generates an answer from $q_{i}$ alone.
Together, intent consistency is ensured, where the reasoning process is always aligned with the purpose of the edit.
4 Experiment
4.1 Experiment Setting
Baselines. We compare CAPE-KG against parameter-updati-ng editors (FT [16], ROME [3], MEMIT [4]) and parameter-preserving method including MeLLo [14], PokcMQA [7], and KeDKG [11]. CAPE-KG is a model-agnostic framework; for fairness, we follow the experimental configuration of KeDKG [11], including detector models and backbones.
Datasets. Experiments are conducted on benchmark MQuAKE  [14], includes MQuAKE-CF-3K and MQuAKE-T used for evaluation. Few-shot question decomposition demos are drawn from MQuAKE-CF after removing overlaps with the evaluation data, and the progressive retrieval parameters ($\tau=0.6$, $\lambda=1$) are selected via grid search on it.
Evaluation Metrics. The framework is evaluated by Match Accuracy (M-Acc), which measures the correctness of the final predicted answer against the benchmark ground truth, and Hop Accuracy (H-Acc), which measures the accuracy of intermediate hops against annotated ground-truth chains.
4.2 Results
Main Results Table 1 shows that CAPE-KG consistently outperforms all baselines across all backbones and datasets. On average, we achieve an absolute improvement of 20.67% in M-Acc and 19.16% in H-Acc on MQuAKE-CF-3K, an absolute improvement of 21.42% in M-Acc and 16.9% in H-Acc on MQuAKE-T. These substantial gains demonstrate the effectiveness of enforcing multiperspective consistency in PPKE for MHQA. The improvements hold across LLaMA-2 [15], Vicuna [17], and GPT-3.5 backbones, indicating that the performance gains in accuracy are model-independent and arise from this consistency-aware architecture.
Consistency Across Batch Edits We observed accuracy drops in both M-Acc and H-Acc on baseline as edit batch size increases, due to overwrites and cross-case contamination. In addition to higher accuracy, CAPE-KG’s multi-layer design with isolated overlays preserves consistency under larger batch sizes, leading to smaller fluctuations. The absolute gap between single-edit and all-edit settings across two datasets averages only 0.65% for M-Acc and 0.54% for H-Acc, further highlighting the improvement in consistency of our approach.
Knowledge Boundary Consistency. CAPE-KG has effectively improved the knowledge boundary consistency through avoiding possible contamination. In the case shown in 2, KEDKG [11] has the potential inconsistency of mixing unrelated edit from other cases, (Black Pink, genre, EDM), into reasoning as fact. EDM is retrieved instead of K-pop, which breaks the reasoning chain, while CAPE-KG maintains knowledge boundary consistency by using this multi-layer design.
Update Consistency. CAPE-KG prevents the possible inconsistency caused by overwriting conflicting edits in a global KG. In the case shown in 2, two edits (K-pop, origin_country, Turkey), (K-pop, origin_country, Germany) cannot coexist under the conflict-solving of Kedkg [11], so one remains while the other is discarded. CAPE-KG stores each edit in a case-specific overlay, preserving both without interference. This design keeps per-case reasoning aligned with the case-isolated edit assumption and ensures update consistency.
Intent Consistency. As shown in Fig. 3, when a subquestion is directly related to an edited fact, KeDKG [11] falls back to LLM answering directly under low-confidence entity extraction. This fallback would query unedited factual knowledge, which is inconsistent with the intent of editing. In contrast, CAPE-KG routes edit-related queries to the overlay layer via the retrieval router, ensuring that edits are explicitly applied in reasoning. Moreover, at the failure stage of progressive retrieval, the LLM is also queried, but with the edit injected into the query context. This guarantees that even when a component of CAPE-KG is uncertain, the reasoning process remains aligned with the intended update, in this way preserving intent consistency.
4.3 Ablation Studies
Table 2 shows the ablation study of CAPE-KG by disenabling its three components: KG Construction, Knowledge Update, and Knowledge Retrieval. Performance improves steadily as more modules are introduced. Among them, the Retrieval module contributes the most significant gains, highlighting the importance of enforcing intent consistency. In contrast, KG Construction and Knowledge Update have relatively minor impacts, as their effectiveness depends on the prevalence of conflicts in the dataset. Nevertheless, the combination of all three modules results in the highest accuracy, demonstrating that multiperspective consistency is best achieved when the full framework is employed, which makes the most effective PPKE on MHQA.
5 Conclusion
In this paper, we propose CAPE-KG, which advances PPKE for MHKE by enforcing knowledge boundary consistency, update consistency, and intent consistency through components of the framework, including KG construction, knowledge update, and knowledge retrieval. In future work, we expect to extend CAPE-KG to continual knowledge editing, ensuring consistency as edits evolve over time.